OpenStudy (anonymous):

The human body can safely stand an acceleration of 9 times that due to gravity. With what minimum radius of curvature may a pilot safely turn the plane upward at the end of a dive if the plane's speed is 770kmph?

OpenStudy (anonymous):

9g=v^2/r g is acceleration due to gravity v=velocity r= radius of curvature

OpenStudy (anonymous):

formula comes by force balancing ma=mv^2/r m=mass a=acceleration

OpenStudy (anonymous):

The confusion comes from what a human body feels! Is it the net force = mV^2/r or Normal force on you = mV^2/r

OpenStudy (anonymous):

its the normal force acting on u which makes the net force.

OpenStudy (anonymous):

yes, but do you feel the net force as 9mg or just the normal force as 10mg

OpenStudy (anonymous):

i think its 9mg only

OpenStudy (anonymous):

as our body cant safely withstand 10mg as per problem,

OpenStudy (anonymous):

and also the net effect is 9mg taking into account the mg force